OpenStudy (hulkart):

Find the arithmetic sum 52 + 59 + 66 + … + 248. S29 = 4,350 S28 = 4,200 S30 = 4,500 S27 = 4,050

Nnesha (nnesha):

do you know the formula to find arithmetic sum ??

OpenStudy (hulkart):

no

Nnesha (nnesha):

formula to find arithmetic sum \[\huge\rm s_n = \frac{ n(a_1+a_n) }{ 2 }\] a_1=first term a_n = last term n = total number of terms because you don't know how many terms are there so first step is to find n to find n apply this formula \[\huge\rm a_n = a_1 + (n-1)d\]

Nnesha (nnesha):

replace variables by their values and solve for n \[\large\rm a_n = a_1 +(n-1)d\] d = common difference do you know how to find common difference from that sequence ?

OpenStudy (jhannybean):

in your sequence, \((\color{red}{52} + 59 + 66 + … + \color{blue}{248})\) your \(\color{blue}{a_n = 248}\) and your \(\color{red}{a_1 = 52}\)

OpenStudy (jhannybean):

To find d, subtract \(a_3 - a_2 = 66-59 =~? ~,~ a_2-a_1 = 59-52=~?\) \[a_n=~?\]\[a_1=~?\]\[d=~?\] Use those two values in the equation @Nnesha provided for you to find \(n\). \(a_n=a+1+(n-1)d\)